OpenStudy (anonymous):

x^3+3x^2+3x+9 Factor the four term polynomial by grouping

OpenStudy (anonymous):

factor out the common terms from x^3+3x^2 and 3x+9

OpenStudy (anonymous):

group the first two terms together and group the last two terms together. then factor each grouping... what you got??

OpenStudy (anonymous):

6 I am not sure

OpenStudy (anonymous):

what's common to x^3+3x^2 ?

OpenStudy (anonymous):

look at it like this if you're unsure x^3+3x^2 is the same as x(x^2) +3(x^2)

OpenStudy (anonymous):

This is my first algebra class so I am confused

OpenStudy (cwrw238):

factor x^3 + 3x^2 common factor is x^2 so we get x^2((x + 3) do you get that ok?

OpenStudy (anonymous):

ok

OpenStudy (cwrw238):

right now lets look at the third and 4th terms + 3x + 9 what is common factor in 3x and 9?

OpenStudy (anonymous):

12

OpenStudy (cwrw238):

no - what common factor means is the number ( or term) that divides exactly into both 3x and 9

OpenStudy (anonymous):

3

OpenStudy (cwrw238):

yes - actually its the Greatest Common factor (GCF) of 3x and 9

OpenStudy (cwrw238):

dividing 3x + 9 by 3 = 3( x + 3)

OpenStudy (cwrw238):

ok?

OpenStudy (anonymous):

ok

OpenStudy (cwrw238):

so bringing them altogether x^3+3x^2+3x+9 = x^2(x + 3) + 3(x + 3) note that (x+3) is common to both parts so we can simplify more

OpenStudy (cwrw238):

if you have something like 3x + 3y factors to 3(x + y) because 3 is common t0 both terms compare this with last post where (x + 3) is common factor

OpenStudy (anonymous):

How would i simplify it

OpenStudy (cwrw238):

i took 3 out in last problem so here you take (x + 3) out

OpenStudy (cwrw238):